OpenStudy (josedavid):

Pedro has created the function f(x) = the quantity of 4x minus 3, divided by 2 to represent the number of assignments he has completed, where x represents the number of weeks in the course. Pedro discovers that, using the inverse function to solve for x = 30, he can predict when he will have 30 assignments completed. Explain to Pedro how to accomplish this, using complete sentences.

OpenStudy (cal.lavender):

30 = (4*x-3)/2 60 = (4*x-3) - multiplied both sides by 2 63 = 4*x - added 3 to both sides. 63/4 = x so after 63/4 weeks which is 15.7 5 weeks or 15 weeks and 3/4 OR 15 weeks, 5 days and 6 hours. Hope I helped :))

OpenStudy (josedavid):

thanks